11.A variable has a mean of 100 and a standard deviation of 24. Sixteen observations of this variable have a mean of 113 and a s
Answer #1

Solution: Given that μ = 100, σ = 24, n = 16

a. Z = (X-μ)/(σ/sqrt(n))
= (113-100)/(24/sqrt(16))
= 2.1667

b. t = (X-μ)/(σ/sqrt(n))
= (113-100)/(16/sqrt(16))
= 3.25